OpenStudy (anonymous):

If a rectangular lot measures 160 ft. x 110 ft., and 15 foot set backs are required from each lot line, how much area is left? If 65 cars require 32,500 square feet, how many cars will the small rectangle hold? There will be ______ square feet in the smaller lot. The number of cars is____. (Use a whole number.)

jhonyy9 (jhonyy9):

your opinion ?

OpenStudy (anonymous):

160x110=17,600

jhonyy9 (jhonyy9):

right and how you need to continue it ?

jhonyy9 (jhonyy9):

so if 15 foot is set backs

OpenStudy (anonymous):

17,600/15=1,173.33

jhonyy9 (jhonyy9):

so if you set backs 25 foot from each lot line how many area will remained from what you have calculated firstly ?

jhonyy9 (jhonyy9):

how do you calcule this ?

jhonyy9 (jhonyy9):

so what mean 15 foot set backs from each lot line ?

jhonyy9 (jhonyy9):

how you calcule how much area was required ?

OpenStudy (anonymous):

17,000x15

OpenStudy (anonymous):

lxw

OpenStudy (anonymous):

I"m lost i'm not sure

OpenStudy (anonymous):

LxW

jhonyy9 (jhonyy9):

so firstly you calculated area of this surface with 160 ft X 110 ft yes ? than now from this you need calcule how much ,how many area is left if this 15 ft was left back

jhonyy9 (jhonyy9):

sorry my internet conection was wrang so can you continue it ?

OpenStudy (anonymous):

17,600/15

jhonyy9 (jhonyy9):

how do you think if from each lot lines are required back 15 foot how you calcule that how many area is left ?

OpenStudy (anonymous):

12in=1foot

OpenStudy (anonymous):

12x15=180

jhonyy9 (jhonyy9):

how you think what mean that 15 foot set backs are required from each lot line ?

OpenStudy (anonymous):

i don't know can you explain me

jhonyy9 (jhonyy9):

so there are 160 ft there what mean ft ?

OpenStudy (anonymous):

160x15=2,400

jhonyy9 (jhonyy9):

why 160 time 15 ?

jhonyy9 (jhonyy9):

@Mikael can you intervening here ?

jhonyy9 (jhonyy9):

@mark_o. do you understand it right ?

OpenStudy (anonymous):

hmm..not quite.. im still figuring out what the prob means..lol

OpenStudy (anonymous):

you go ahead we are listening

jhonyy9 (jhonyy9):

how do you think what mean 15 foot set backs are required from each lot line ?

jhonyy9 (jhonyy9):

so than how much area was left ?

OpenStudy (anonymous):

so is that mean 160 ft divided by 15 ft? =10.66 ft

OpenStudy (anonymous):

go ahead @ jhonyy9 continue and we are listening

OpenStudy (anonymous):

hmm i think 15 ft set back means take awy 15ft at each end of 160 ft that will be 160-30=130ft..... and at 110-30=80 so the remaining rectangle has 80x130==10,400 sq ft ..ans... then it is required that 32,500 sq ft/65 cars=500 therefore 10,400/500 =20.8 cars use 20 car as approx ...ans

jhonyy9 (jhonyy9):

but 160 ft is a measure og the lenght of and the width is 110 ft

jhonyy9 (jhonyy9):

og wann being of ---sorry

jhonyy9 (jhonyy9):

so if 15 foot is set back so than how much area was required from the first calculated area ?

OpenStudy (anonymous):

so i think it is 80x130=10,400 sq ft

jhonyy9 (jhonyy9):

@mark_o. how do you think not is very very interesting that there are ft and foot too so how can you calcule with two unit of measure ? this is possibile ?

jhonyy9 (jhonyy9):

so ft mean feet and foot not is equal feet yes ?

OpenStudy (anonymous):

hmm ft = feet= foot are the same

jhonyy9 (jhonyy9):

but there the asker have wrote wrong ???

OpenStudy (anonymous):

not really a foot is singular and feet are plural

jhonyy9 (jhonyy9):

ok so than how do you think if there have wrote that 15 foot are required back from each line so than what mean this again ?

OpenStudy (anonymous):

i think they want 15 foot set backs from each line is like having a border line at 160 ft and at 110 ft

jhonyy9 (jhonyy9):

ok i think the same

OpenStudy (anonymous):